Passing the ball between two players during a soccer game is a critical skill for the success of a team. A random sample of passes made by two teams in the 2010 World Cup was​ drawn, and the number of successful passes in each sample was counted. Construct a

9999​%

confidence interval to estimate the difference in accuracy between the two​ countries' passes. What conclusions can be​ made?Passing the ball between two players during a soccer game is a critical skill for the success of a team. A random sample of pCountry 1 Xq = 67 n = 74 Country 2 X2 = 74 n2 = 88
Answer #1

criven, n=67 ny = 74 ni = 74 n2 = 88. - - * = 0.905 Part = 0.641 Po= mit me 57+ 14 = 0.87 Olt too 0 . 8 ntna 74+88 %= l-Po =

- 0.064 + 0.141 > C -0.077 0.205). what conclusion can be made D Because this confidence interval does not include zero, ther